OpenStudy (anonymous):

Simplify to exponential form. 8 divided by 7sqrt(x^15)

OpenStudy (abb0t):

take the natural log?

OpenStudy (anonymous):

dont know how.. i missed a bunch of the school and i dont know how to do any of it

OpenStudy (abb0t):

I'm assuming the problem looks like this: \[\frac{ 8 }{ 7\sqrt{x^{15}} }\]

OpenStudy (anonymous):

yes.

OpenStudy (abb0t):

\[\frac{ 8 }{ 7 }(\frac{ 1 }{ (x^{15})^\frac{ 1 }{ 2 } } ) \] \[\frac{ 8 }{ 7 }[\ln(1)-\frac{ 1 }{ 2 }\ln(x^{15})]\]

OpenStudy (anonymous):

... i have no idea.. what that is saying.. my options are.. a.8x -7/15 b.8x 15/7 c.8x -15/7 d.8x 7/15